OpenStudy (anonymous):

Solve 2 – 5x = x – 10 x = –3 x = –2 x = 2 x = 3

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

Or A?

OpenStudy (anonymous):

so @Deldrickg you want to isolate the varible, this means that you want all the x's on the same side. 2-5x=x-10 so then add 10 to both sides and add 5x to both sides 12=6x now divide both sides by 6 / formulas 2=x

OpenStudy (anonymous):

C

OpenStudy (anonymous):

thanku

mathslover (mathslover):

ok let me explain u this clearly : \[\huge{2-5x=x-10}\] \[\huge{2-5x-x=-10}\] \[\huge{2-6x=-10}\] \[\huge{2+10=6x}\] \[\huge{12=6x}\] \[\huge{x=\frac{12}{6}=2}\]

OpenStudy (anonymous):

you also deserve one @mathslover